The density of aluminum is 2.70 g/mL. If you use 108 g of aluminum foil in an experiment, what is the volume of the aluminum in milliliters?
1
Expert's answer
2020-12-10T07:44:43-0500

p=m/V.

V=m/p = 108/2.7 = 40 mL.